# Quantum algorithm for partial differential equations of non-conservative systems with spatially varying parameters

### v1

@@ III.3 Coefficient oracle by matrix product state
$K_{a_{n_{\mathrm{a}}}b_{n_{\mathrm{a}}}}^{[n_{\mathrm{a}}]}$ =$=\begin{cases}-a_{n_{\mathrm{a}}}2^{n_{\mathrm{a}}-1-n_{\mathrm{frac}}}&\text{ if }b_{n_{\mathrm{a}}}=n_{\mathrm{a}}\\ 1&\text{otherwise}\end{cases}$ (29)
@@
That is, $\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}k_{a}\ket{a}$ can be exactly represented as an MPS with the bond dimension of $n_{\mathrm{a}}$. Next, we derive an MPS that represents $\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}(1+k_{a}^{2})\ket{a}$ from an MPS representing $\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}k_{a}\ket{a}$ as follows:
$\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}(1+k_{a}^{2})\ket{a}$ $=\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\ket{a}+\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}k_{a}^{2}\ket{a}$
$=\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\ket{a}$
$\quad+\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\sum_{b_{n_{\mathrm{a}}},\dots b_{1}=1}^{n_{\mathrm{a}}}\sum_{b^{\prime}_{n_{\mathrm{a}}},b^{\prime}_{1}=1}^{n_{\mathrm{a}}}K_{a_{n_{\mathrm{a}}}b_{n_{\mathrm{a}}}}^{[n_{\mathrm{a}}]}\dots K_{b_{2}a_{1}}^{[1]}K_{a_{n_{\mathrm{a}}}b^{\prime}_{n_{\mathrm{a}}}}^{[n_{\mathrm{a}}]}\dots K_{b^{\prime}_{2}a_{1}}^{[1]}\ket{a_{n_{\mathrm{a}}}a_{n_{\mathrm{a}}-1}\dots a_{1}}$
$=\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\sum_{b_{n_{\mathrm{a}}},b_{1}=1}^{n_{\mathrm{a}}^{2}+1}S_{a_{n_{\mathrm{a}}}b_{n_{\mathrm{a}}}}^{[n_{\mathrm{a}}]}S_{b_{n_{\mathrm{a}}}a_{n_{\mathrm{a}}-1}b_{n_{\mathrm{a}}-1}}^{[n_{\mathrm{a}}-1]}\dots S_{b_{3}a_{2}b_{2}}^{[2]}S_{b_{2}a_{1}}^{[1]}\ket{a_{n_{\mathrm{a}}}a_{n_{\mathrm{a}}-1}\dots a_{1}},$ (32)
@@
$S_{a_{n_{\mathrm{a}}}b_{n_{\mathrm{a}}}}^{[n_{\mathrm{a}}]}$ =$=\begin{cases}K_{a_{n_{\mathrm{a}}}\lfloor b_{n_{\mathrm{a}}}/n_{\mathrm{a}}\rfloor}^{[n_{\mathrm{a}}]}K_{a_{n_{\mathrm{a}}}\texttt{mod}(b_{n_{\mathrm{a}}},n_{\mathrm{a}})}^{[n_{\mathrm{a}}]}&\text{ if }b_{n_{\mathrm{a}}}\leq n_{\mathrm{a}}^{2}\\ 1&\text{otherwise}\end{cases}$ (33)
$S_{b_{m+1}a_{m}b_{m}}^{[m]}$ =$=\begin{cases}K_{\lfloor b_{m+1}/n_{\mathrm{a}}\rfloor a_{m}\lfloor b_{m}/n_{\mathrm{a}}\rfloor}^{[m]}K_{\texttt{mod}(b_{m+1},n_{\mathrm{a}})a_{m}\texttt{mod}(b_{m},n_{\mathrm{a}})}^{[m]}&\text{ if }b_{m}\leq n_{\mathrm{a}}^{2}\\ \delta_{b_{m+1}b_{m}}&\text{otherwise}\end{cases}$ (34)
$S_{b_{2}a_{1}}^{[1]}$ =$=\begin{cases}K_{\lfloor b_{2}/n_{\mathrm{a}}\rfloor a_{1}}^{[1]}K_{\texttt{mod}(b_{2},n_{\mathrm{a}})a_{1}}^{[1]}&\text{ if }b_{2}\leq n_{\mathrm{a}}^{2}\\ 1&\text{otherwise}\end{cases},$ (35)
That is, $\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}(1+k_{a}^{2})\ket{a}$ can also be exactly represented by an MPS with the bond dimension of $n_{\mathrm{a}}^{2}+1$. However, we were unable to find an analytical method to represent the target $\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\sqrt{1+k_{a}^{2}}\ket{a}$ in an MPS form. Therefore, we decided to obtain the target MPS through optimization, which utilizes $\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}(1+k_{a}^{2})\ket{a}$. Consider the following vector-valued function:
$\mathcal{F}(\bm{\Psi}):=\texttt{diag}(\bm{\Psi})\bm{\Psi}-\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}(1+k_{a}^{2})\ket{a},$ (36)
where diag is an operator that returns a diagonal matrix whose diagonal components corresponds to the input vector $\bm{\Psi}$. Obviously, $\mathcal{F}(\bm{\Psi})={\bf 0}$ iff $\bm{\Psi}=\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\sqrt{1+k_{a}^{2}}\ket{a}$, which is our target state. In this study, we implement the Newton’s method to find $\bm{\Psi}$ such that $\mathcal{F}(\bm{\Psi})={\bf 0}$. That is, using the Taylor expansion:
@@
Eq. (39) is a system of linear equations represented in MPS, and in this study, it was solved using the modified alternating least square (MALS) method [43]. Through this procedure, we can obtain an MPS that approximates $\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\sqrt{1+k_{a}^{2}}\ket{a}$ in the MPS form. Finally, by solving the system of linear equations:
$\texttt{diag}(\bm{\Psi})\bm{\Phi}=\sqrt{\frac{2^{n_{\mathrm{frac}}}}{\pi}}\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\ket{a},$ (40)
we obtain the MPS $\bm{\Phi}$ that approximates the target state $\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\sqrt{c_{a}}\ket{a}$ with $c_{a}=2^{-n_{\mathrm{frac}}}/\pi(1+k_{a}^{2})$. We used the MALS method to solve Eq. (40). Through these procedures, if the bond dimension of the MPS is kept at $\mathcal{O}(\mathrm{poly}(n_{\mathrm{a}}))$, we can conduct a normalization of $\bm{\Phi}$ and construct an MPS that approximates $(1/\sqrt{\|\bm{c}\|_{1}})\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\sqrt{c_{a}}\ket{a}$, within a time complexity of $\mathcal{O}(\mathrm{poly}(n_{\mathrm{a}}))$. Specifically, let $r_{\Psi}$ and $r_{\Phi}$ denote the bond dimensions for $\bm{\Psi}$ and $\bm{\Phi}$, respectively. Since the time complexity of the MALS is $\mathcal{O}(\max(r_{x},r_{b})^{3}r_{A}^{2}n)$ for solving a linear system $Ax=b$ where $A$, $x$ and $b$ are MPO and MPSs with the bond dimensions of $r_{A}$, $r_{x}$ and $r_{b}$, respectively [43], the time complexity for solving Eq. (39) is $\mathcal{O}((r_{\Psi}^{2}+n_{\mathrm{a}}^{2}+1)^{3}r_{\Psi}^{2}n_{\mathrm{a}})$. Also, the time complexity for solving Eq. (40) is $\mathcal{O}(r_{\Phi}^{3}r_{\Psi}^{2}n_{\mathrm{a}})$.
The above procedures are conducted entirely through classical calculations of tensor networks. To realize this classical state on a quantum device, or in other words to obtain the coefficient oracle $O_{\mathrm{coef}}$, we use the scheme for converting an MPS into a quantum circuit [41, 42]. Note that no access to the full state vector with the dimension of $2^{n_{\mathrm{a}}}$ is necessary. The specific procedure is summarized in Algorithm 2.
@@ III.4 Simulation algorithm and its complexity
    Prepare the system register consisting of $n$ qubits to embed the solution $\bm{w}(t)$ of the ODE and the ancilla register consisting of $n_{\mathrm{a}}$ ancilla qubits for the LCU. The required number of integration points are $\mathcal{O}(\|\bm{L}\|_{2}T/\varepsilon^{2})$ where $T$ is a simulation time and $\varepsilon$ is a truncation error of integration [25], which results in the number of ancilla qubits, $\mathcal{O}(\log(\|\bm{L}\|_{2}T/\varepsilon^{2}))$.
@@
    Construct the quantum circuit for the coefficient oracle $O_{\mathrm{coef}}$ on $n_{\mathrm{a}}$-qubit ancilla system via the tensor network technique as described in Section III.3. The time complexity for solving Eq. (39) is $\mathcal{O}((r_{\Psi}^{2}+n_{\mathrm{a}}^{2}+1)^{3}r_{\Psi}^{2}n_{\mathrm{a}})$ while the time complexity for solving Eq. (40) is $\mathcal{O}(r_{\Phi}^{3}r_{\Psi}^{2}n_{\mathrm{a}})$.
@@
    Initialize all the qubits to |$\ket{0}^{\otimes n_{\mathrm{a}}}\otimes\ket{0}^{\otimes n}$.
@@
    Apply the coefficient oracle for the LCU, $O_{\mathrm{coef}}:\ket{0}\rightarrow(1/\sqrt{\|\bm{c}\|_{1}})\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\sqrt{c_{a}}\ket{a}$, to the ancilla register.
@@
    Alternatively apply the Hamiltonian simulation oracle $O_{H}$ to the system register and the select oracle $\texttt{SEL}_{L}$ to the whole system $r$ times, where $r:=T/\tau$ is the number of steps for time evolution. Since the Hamiltonians $\bm{L}$ and $\bm{H}$ are represented by the operators $\sigma_{01}$, $\sigma_{10}$, $\sigma_{00}$, $\sigma_{11}$ and $I$, the time evolution operators of them, $O_{L}(\tau):=\exp(-i\bm{L}\tau)$ and $O_{H}(\tau):=\exp(-i\bm{H}\tau)$, can be implemented efficiently [29]. The select oracle $\texttt{SEL}_{L}(\tau):=\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\ket{a}\!\bra{a}\otimes e^{-ik_{a}\bm{L}\tau}$ is implemented as
    $\texttt{SEL}_{L}(\tau)$ $:=\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\ket{a}\!\bra{a}\otimes e^{-ik_{a}\bm{L}\tau}$
    $=\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\ket{a}\!\bra{a}\otimes\exp\left(-i\left(-a_{n_{\mathrm{a}}-1}2^{n_{\mathrm{a}}-1}+\sum_{m=0}^{n_{\mathrm{a}}-2}a_{m}2^{m}\right)2^{-n_{\mathrm{frac}}}\bm{L}\tau\right)$
    $=\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\ket{a}\!\bra{a}\otimes\exp\left(ia_{n_{\mathrm{a}}-1}2^{n_{\mathrm{a}}-1-n_{\mathrm{frac}}}\bm{L}\tau\right)\prod_{m=0}^{n_{\mathrm{a}}-2}\exp\left(-ia_{m}2^{m-n_{\mathrm{frac}}}\bm{L}\tau\right)$
    $=\left(I^{\otimes n_{\mathrm{a}}}+\ket{1}\!\bra{1}_{n_{\mathrm{a}}-1}\otimes O_{L}(-2^{n_{\mathrm{a}}-1-n_{\mathrm{frac}}}\tau)\right)\prod_{m=0}^{n_{\mathrm{a}}-2}\left(I^{\otimes n_{\mathrm{a}}}+\ket{1}\!\bra{1}_{m}\otimes O_{L}(2^{m-n_{\mathrm{frac}}}\tau)\right),$ (41)
@@
    Apply the inverse of the coefficient oracle, $O_{\mathrm{coef}}^{\dagger}$, to the ancilla register. This procedure prepares a quantum state approximating $(1/\|\bm{c}\|_{1})\ket{0}^{\otimes n_{\mathrm{a}}}\otimes\exp\left(-\bm{A}t\right)\ket{w(0)}+\ket{\perp}$ in the sense of numerical integration, where $\ket{\perp}$ represents a state orthogonal to the first term.

### v2

@@ III.3 Coefficient oracle realized using matrix product state
$K_{a_{n_{\mathrm{anc}}}b_{n_{\mathrm{anc}}}}^{[n_{\mathrm{anc}}]}$ =$=\begin{cases}-a_{n_{\mathrm{anc}}}2^{n_{\mathrm{anc}}-1-n_{\mathrm{frac}}}&\text{ if }b_{n_{\mathrm{anc}}}=n_{\mathrm{anc}}\\ 1&\text{otherwise}\end{cases}$ (31)
@@
That is, $\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}k_{a}\ket{a}$ can be exactly represented as an MPS with the bond dimension of $n_{\mathrm{anc}}$. Note that it is enough to store matrices in Eqs. (31)–(33) to represent the MPS and we do not have to compute the exponentially many summation in Eq. (30). Next, we derive an MPS that represents $\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}(1+k_{a}^{2})\ket{a}$ from an MPS representing $\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}k_{a}\ket{a}$ as follows:
$\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}(1+k_{a}^{2})\ket{a}$ $=\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\ket{a}+\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}k_{a}^{2}\ket{a}$
$=\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\ket{a}$
$\quad+\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\sum_{b_{n_{\mathrm{anc}}},\dots b_{1}=1}^{n_{\mathrm{anc}}}\sum_{b^{\prime}_{n_{\mathrm{anc}}},b^{\prime}_{1}=1}^{n_{\mathrm{anc}}}K_{a_{n_{\mathrm{anc}}}b_{n_{\mathrm{anc}}}}^{[n_{\mathrm{anc}}]}\dots K_{b_{2}a_{1}}^{[1]}K_{a_{n_{\mathrm{anc}}}b^{\prime}_{n_{\mathrm{anc}}}}^{[n_{\mathrm{anc}}]}\dots K_{b^{\prime}_{2}a_{1}}^{[1]}\ket{a_{n_{\mathrm{anc}}}a_{n_{\mathrm{anc}}-1}\dots a_{1}}$
$=\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\sum_{b_{n_{\mathrm{anc}}},b_{1}=1}^{n_{\mathrm{anc}}^{2}+1}S_{a_{n_{\mathrm{anc}}}b_{n_{\mathrm{anc}}}}^{[n_{\mathrm{anc}}]}S_{b_{n_{\mathrm{anc}}}a_{n_{\mathrm{anc}}-1}b_{n_{\mathrm{anc}}-1}}^{[n_{\mathrm{anc}}-1]}\dots S_{b_{3}a_{2}b_{2}}^{[2]}S_{b_{2}a_{1}}^{[1]}\ket{a_{n_{\mathrm{anc}}}a_{n_{\mathrm{anc}}-1}\dots a_{1}},$ (34)
@@
$S_{a_{n_{\mathrm{anc}}}b_{n_{\mathrm{anc}}}}^{[n_{\mathrm{anc}}]}$ =$=\begin{cases}K_{a_{n_{\mathrm{anc}}}\lfloor b_{n_{\mathrm{anc}}}/n_{\mathrm{anc}}\rfloor}^{[n_{\mathrm{anc}}]}K_{a_{n_{\mathrm{anc}}}\texttt{mod}(b_{n_{\mathrm{anc}}},n_{\mathrm{anc}})}^{[n_{\mathrm{anc}}]}&\text{ if }b_{n_{\mathrm{anc}}}\leq n_{\mathrm{anc}}^{2}\\ 1&\text{otherwise}\end{cases}$ (35)
$S_{b_{m+1}a_{m}b_{m}}^{[m]}$ =$=\begin{cases}K_{\lfloor b_{m+1}/n_{\mathrm{anc}}\rfloor a_{m}\lfloor b_{m}/n_{\mathrm{anc}}\rfloor}^{[m]}K_{\texttt{mod}(b_{m+1},n_{\mathrm{anc}})a_{m}\texttt{mod}(b_{m},n_{\mathrm{anc}})}^{[m]}&\text{ if }b_{m}\leq n_{\mathrm{anc}}^{2}\\ \delta_{b_{m+1}b_{m}}&\text{otherwise}\end{cases}$ (36)
$S_{b_{2}a_{1}}^{[1]}$ =$=\begin{cases}K_{\lfloor b_{2}/n_{\mathrm{anc}}\rfloor a_{1}}^{[1]}K_{\texttt{mod}(b_{2},n_{\mathrm{anc}})a_{1}}^{[1]}&\text{ if }b_{2}\leq n_{\mathrm{anc}}^{2}\\ 1&\text{otherwise}\end{cases}.$ (37)
That is, $\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}(1+k_{a}^{2})\ket{a}$ can also be exactly represented by an MPS with the bond dimension of $n_{\mathrm{anc}}^{2}+1$. Note again that the calculation of Eq. (III.3) can be performed via calculation of each matrix in Eqs. (35)–(37) and we do not have to compute the exponentially many summation in Eq. (III.3). However, we were unable to find an analytical method to represent the target $\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\sqrt{1+k_{a}^{2}}\ket{a}$ in an MPS form. Therefore, we decided to obtain the target MPS through optimization, which utilizes $\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}(1+k_{a}^{2})\ket{a}$. Consider the following vector-valued function:
$\mathcal{F}(\bm{\Psi}):=\texttt{diag}(\bm{\Psi})\bm{\Psi}-\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}(1+k_{a}^{2})\ket{a},$ (38)
where diag is an operator that returns a diagonal matrix whose diagonal components corresponds to the input vector $\bm{\Psi}$. Obviously, $\mathcal{F}(\bm{\Psi})={\bf 0}$ iff $\bm{\Psi}=\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\sqrt{1+k_{a}^{2}}\ket{a}$, which is our target state. In this study, we implement the Newton’s method to find $\bm{\Psi}$ such that $\mathcal{F}(\bm{\Psi})={\bf 0}$. That is, using the Taylor expansion:
@@
Eq. (41) is a system of linear equations represented in MPS, and in this study, it was solved using the modified alternating least square (MALS) method [48]. Through this procedure, we can obtain an MPS that approximates $\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\sqrt{1+k_{a}^{2}}\ket{a}$ in the MPS form. Finally, by solving the system of linear equations:
$\texttt{diag}(\bm{\Psi})\bm{\Phi}=\sqrt{\frac{2^{n_{\mathrm{frac}}}}{\pi}}\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\ket{a},$ (42)
we obtain the MPS $\bm{\Phi}$ that approximates the target state $\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\sqrt{c_{a}}\ket{a}$ with $c_{a}=2^{-n_{\mathrm{frac}}}/\pi(1+k_{a}^{2})$. We used the MALS method to solve Eq. (42). Through these procedures, if the bond dimension of the MPS is kept at $\mathcal{O}(\mathrm{poly}(n_{\mathrm{anc}}))$, we can conduct a normalization of $\bm{\Phi}$ and construct an MPS that approximates $(1/\sqrt{\|\bm{c}\|_{1}})\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\sqrt{c_{a}}\ket{a}$, within a time complexity of $\mathcal{O}(\mathrm{poly}(n_{\mathrm{anc}}))$. Specifically, let $r_{\Psi}$ and $r_{\Phi}$ denote the bond dimensions for $\bm{\Psi}$ and $\bm{\Phi}$, respectively. Since the time complexity of the MALS is $\mathcal{O}(\max(r_{x},r_{b})^{3}r_{A}^{2}n)$ for solving a linear system $Ax=b$ where $A$, $x$ and $b$ are MPO and MPSs with the bond dimensions of $r_{A}$, $r_{x}$ and $r_{b}$, respectively [48], the time complexity for solving Eq. (41) is $\mathcal{O}((r_{\Psi}^{2}+n_{\mathrm{anc}}^{2}+1)^{3}r_{\Psi}^{2}n_{\mathrm{anc}})$. Also, the time complexity for solving Eq. (42) is $\mathcal{O}(r_{\Phi}^{3}r_{\Psi}^{2}n_{\mathrm{anc}})$.
Figure 3: Constructing a quantum circuit from an MPS with the bond dimension of 2.
The above procedures are conducted entirely through classical calculations of tensor networks. To realize this classical state on a quantum device, or in other words to obtain the coefficient oracle $O_{\mathrm{coef}}$, we use the scheme for converting an MPS into a quantum circuit [44, 46]. Figure 3 illustrates the procedure of constructing a quantum circuit from an MPS with the bond dimension of 2 for 5 qubits based on Ref. [44]. Let $\bm{\Phi}\approx\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\sqrt{c_{a}}\ket{a}$ with $c_{a}=2^{-n_{\mathrm{frac}}}/\pi(1+k_{a}^{2})$ be a right canonical MPS. For the rightmost matrix (topmost matrix in the figure), $\Psi^{[1]}$, which is unitary, we set $U^{[1]}=\Psi^{[1]}$. For $51<m<n_{a}=$ 5, we set the $4\times 4$ unitary $U^{[m]}$ as $U_{2a_{m}+b_{m},2b_{m+1}}^{[m]}=\Psi_{b_{m+1}a_{m}b_{m}}^{[m]}$ for even-numbered column and we set the odd-numbered column $U_{2a_{m}+b_{m},2b_{m+1}+1}^{[m]}$ so that $U^{[m]}$ can be unitary. For the leftmost matrix, we set the $4\times 4$ unitary $U^{[5]}$ as $U_{2a_{5}+b_{5},0}^{[m]}=\Psi_{a_{5}b_{5}}^{[5]}$ for first column and we set other columns so that $U^{[5]}$ can be unitary. Thus, we can construct a quantum circuit as shown in Fig. 3 from an MPS with the bond dimension of 2. For MPSs with larger bond dimensions, we can also approximately construct the corresponding quantum circuits by repeatedly disentangling the MPS using an MPS with the bond dimension of 2 [44, 46]. Note that no access to the full state vector with the dimension of $2^{n_{\mathrm{anc}}}$ is necessary. The specific procedure is summarized in Algorithm 2.
@@ III.4 Simulation algorithm and its complexity
    Prepare the system register consisting of $n$ qubits to embed the solution $\bm{w}(t)$ of the ODE and the ancilla register consisting of $n_{\mathrm{anc}}$ ancilla qubits for the LCU. The required number of integration points are $\mathcal{O}(\|\bm{L}\|_{2}T/\varepsilon^{2})$ where $T$ is a simulation time and $\varepsilon$ is a truncation error of integration [26], which results in the number of ancilla qubits, $\mathcal{O}(\log(\|\bm{L}\|_{2}T/\varepsilon^{2}))$.
@@
    Construct the quantum circuit for the coefficient oracle $O_{\mathrm{coef}}$ on $n_{\mathrm{anc}}$-qubit ancilla system via the tensor network technique as described in Section III.3. The time complexity for solving Eq. (41) is $\mathcal{O}((r_{\Psi}^{2}+n_{\mathrm{anc}}^{2}+1)^{3}r_{\Psi}^{2}n_{\mathrm{anc}})$ while the time complexity for solving Eq. (42) is $\mathcal{O}(r_{\Phi}^{3}r_{\Psi}^{2}n_{\mathrm{anc}})$.
@@
    Initialize all the qubits to |$\ket{0}^{\otimes n_{\mathrm{anc}}}\otimes\ket{0}^{\otimes n}$.
@@
    Apply the coefficient oracle for the LCU, $O_{\mathrm{coef}}:\ket{0}\rightarrow(1/\sqrt{\|\bm{c}\|_{1}})\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\sqrt{c_{a}}\ket{a}$, to the ancilla register.
@@
    Alternatively apply the Hamiltonian simulation oracle $O_{H}$ to the system register and the select oracle $\texttt{SEL}_{L}$ to the whole system $r$ times, where $r:=T/\tau$ is the number of steps for time evolution. Since the Hamiltonians $\bm{L}$ and $\bm{H}$ are represented by the operators $\sigma_{01}$, $\sigma_{10}$, $\sigma_{00}$, $\sigma_{11}$ and $I$, the time evolution operators of them, $O_{L}(\tau):=\exp(-i\bm{L}\tau)$ and $O_{H}(\tau):=\exp(-i\bm{H}\tau)$, can be implemented efficiently [30]. The select oracle $\texttt{SEL}_{L}(\tau):=\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\ket{a}\!\bra{a}\otimes e^{-ik_{a}\bm{L}\tau}$ is implemented as
    $\texttt{SEL}_{L}(\tau)$ $:=\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\ket{a}\!\bra{a}\otimes e^{-ik_{a}\bm{L}\tau}$
    $=\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\ket{a}\!\bra{a}\otimes\exp\left(-i\left(-a_{n_{\mathrm{anc}}-1}2^{n_{\mathrm{anc}}-1}+\sum_{m=0}^{n_{\mathrm{anc}}-2}a_{m}2^{m}\right)2^{-n_{\mathrm{frac}}}\bm{L}\tau\right)$
    $=\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\ket{a}\!\bra{a}\otimes\exp\left(ia_{n_{\mathrm{anc}}-1}2^{n_{\mathrm{anc}}-1-n_{\mathrm{frac}}}\bm{L}\tau\right)\prod_{m=0}^{n_{\mathrm{anc}}-2}\exp\left(-ia_{m}2^{m-n_{\mathrm{frac}}}\bm{L}\tau\right)$
    =(|$=\left(\ket{0}\!\bra{0}_{n_{\mathrm{anc}}-1}\otimes I^{\otimes n}+\ket{1}\!\bra{1}_{n_{\mathrm{anc}}-1}\otimes O_{L}(-2^{n_{\mathrm{anc}}-1-n_{\mathrm{frac}}}\tau)\right)$
    $\hskip 28.45274pt\prod_{m=0}^{n_{\mathrm{anc}}-2}\left(\ket{0}\!\bra{0}_{m}\otimes I^{\otimes n}+\ket{1}\!\bra{1}_{m}\otimes O_{L}(2^{m-n_{\mathrm{frac}}}\tau)\right),$ (43)
@@
    Apply the inverse of the coefficient oracle, $O_{\mathrm{coef}}^{\dagger}$, to the ancilla register. This procedure prepares a quantum state approximating $(1/\|\bm{c}\|_{1})\ket{0}^{\otimes n_{\mathrm{anc}}}\otimes\exp\left(-\bm{A}t\right)\ket{w(0)}+\ket{\perp}$ in the sense of numerical integration, where $\ket{\perp}$ represents a state orthogonal to the first term.